# Bootstrap 3D Reconstructed Scenes from 3D Gaussian Splatting

### v1

@@ 2.1 Diffusion Models
where $\epsilon_{\theta}$ is a learned noise prediction function, and $\epsilon$ is a random normal. For more details, please see in [27, 28, 29]
@@
Image-to-image generation represents a practical application of diffusion models. This paper specifically addresses a unique form of generation wherein an input image is regenerated by the diffusion model, utilizing a defined broken strength $s_{r}$ to facilitate controlled generation. The magnitude of $s_{r}$ directly influences the degree of deviation in the output image from the original, with larger $s_{r}$ values resulting in greater alterations. Given an input image $I$, upon encoding into the latent space $\mathbf{x}^{I}$, we consider $\mathbf{x}^{I}$ is in one of the $1:T$ time steps based on $s_{r}$. For example, if the $s_{r}=0.1$ and $T=1000$, then during the reverse diffusion process, diffusion is directly applied to $p_{\theta}(\mathbf{x}^{I})$ at the time step $T\times s_{r}=100$. This implies that $p_{\theta}(\mathbf{x}^{I})$ is equated with $p_{\theta}(\mathbf{x}_{100})$ in Eq. 1. Consequently, only the last 100 time steps of the reverse diffusion process are executed to conduct partial regeneration.
@@ 3 Methods
3.1 Boostrapping by Diffusion
In the optimization process of 3D-GS, starting from a sparse set of COLMAP [19, 20] points, 3D-GS model will end up with millions of small and dense 3D Gaussians, where 3D Gaussian points undergo splitting or cloning to produce additional points as detailed in original paper [1]. Since the maximum point size (split) and gradient’s accumulation (clone) threshold are hyperparameters to adjust, we mainly focus on the gradient direction that significantly influences where and how new points will be generated.
For a novel-view image $I$ synthesized by the trained 3D-GS model, it is understood that $I$ may not accurately reflect the ground truth image, due to the presence of unseen parts beyond the training dataset. In such cases, if we treat $I$ as a ”partially degenerated version” of $I^{\prime}$, the fine-grained image that aligns the ground truth more, we can employ a diffusion model to achieve image-to-image integration on $I$ to get the image density $I^{\prime}$ as described in Sec. 2.1. Then $I^{\prime}$ can be represented as a collection of noise added to the synthesized image $I$ (we write in this form just for simplicity and better comprehension, for detail please see [27, 28]):
$I^{\prime}=\sum_{t\in T_{s_{r}}}{{\boldsymbol{\epsilon}}_{\theta}(\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{t}(I,{\boldsymbol{\epsilon}})+\sqrt{1-\bar{\alpha}_{t}}{\boldsymbol{\epsilon}},t)}+I,$ (7)
@@ 3.1 Boostrapping by Diffusion
Under this framework, the degree to which $I^{\prime}$ aligns with the ground truth directly correlates to the improvement in our results. However, achieving this ideal situation in practice is very difficult. This is not only because 1. the currently available open-source diffusion models cannot perform well in generating partial images in all scenarios, but also because 2. many scenarios themselves are complex and unique. Moreover, 3. the very process of bootstrapping introduces a level of uncertainty. Ideally, the constructed 3D scene would appear consistent across all perspectives. However, the variability introduced by the diffusion model makes it challenging to guarantee that scenes regenerated by the diffusion model from one new viewpoint will seamlessly align with those from another. Additionally, 4. not all parts of the synthesized novel-view images should be regenerated, if we apply the diffusion process to the whole image, some already well-trained parts will be distorted to some extent. Therefore, based on these problems, we have explored a more universally usable and more stable method.
3.2 Gradient Direction and Diffusion
Bootstrapping.
Let’s start with the timing of the introduction of the bootstrapping technique. Regarding the issue of overall scene consistency, it requires the start of bootstrapping from a relatively complete stage on reconstructed 3D scenes with small diffusion broken strength. For regenerated novel-view images by diffusion model, if the surrounding context is similar and homogeneous, we can consider that the regenerated content in this context tends to have little variation. The extent of differences $\sum_{t\in T_{s_{r}}}{{\boldsymbol{\epsilon}}_{\theta}(\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{t}(I,{\boldsymbol{\epsilon}})+\sqrt{1-\bar{\alpha}_{t}}{\boldsymbol{\epsilon}},t)}$ part are limited, not solely due to the similarity of the initial context $I$ part, but also because the diffusion process is applied only in the final few timesteps (the closer $t$ is to 0, the smaller the noise introduced at $t$ is [6]).
@@ 3.2 Gradient Direction and Diffusion
After the 3D-GS model can construct a relatively complete scene, bootstrapping is applied to every training image. To get a novel-view image $I^{t}_{n}$ from a training image $I_{t}$, a set of new camera parameters for rendering is acquired. In the COLMAP [19, 20] form, we solely slightly change a training camera’s $qvec$ and $tvec$ to obtain new view cameras, detailed in the appendix. Then we regenerate $I^{t}_{n}$ by diffusion model $D$ to get $I^{t}_{d}$, where $I^{t}_{d}=D(I^{t}_{n})$. In practice, we generate multiple novel views from each training camera.
Controllable Diffusion Direction.
@@
Regenerated content can be categorized into two distinct parts: one that contains a faithful representation of the scene but changed with alterations, and another that exhibits an incomplete or inaccurate depiction of the scene, having been modified by the diffusion model. By managing the ratio of the loss associated with the original training component and that of the bootstrapping component, we can achieve a globally stable training process for the 3D-GS model and attain superior results.
@@
We set our bootstrapping loss term as $\mathcal{L}_{1}$ loss $\mathcal{L}_{b}$ of the novel-view rendered image $I^{t}_{n}$ with the regenerated image $I^{t}_{d}$, where $\mathcal{L}_{b}=\lVert I^{t}_{n}-I^{t}_{d}\lVert$. During training, if bootstrapping is used, then for every training camera, we use a hybrid loss with a small $\lambda_{\text{boot}}$ that
@@
where $\mathcal{L}_{o}$ is the original 3D-GS training loss with ground-truth image [1], $\sum_{i\in N}{\mathcal{L}^{i}_{b}}$ is the sum of the bootstrapping loss introduced by all novel-view camera variants based on that training camera together with its surrounding cameras, and $N$ is the total variant number. For example, if the training camera has both sides of surrounding cameras and we bootstrap each camera with 2 of its variants, then the $N$ is 6. This design choice is motivated by the common characteristic of most 3D reconstruction datasets [83, 84, 85], where cameras positioned in close proximity typically capture adjacent scenes.
@@
When the $\lambda_{\text{boot}}$ is sufficiently small, such as 0.05, it becomes evident that the bootstrapping loss term will not significantly affect the first scenario, in which well-trained parts are subject to alteration. And for those parts or viewing angles that are invisible in training, where their renderings are obviously disordered or lacking in detail, the bootstrapping term is sufficient to make changes that are consistent across multiple viewpoints. Since in our approach, when bootstrapping scenarios involve the same parts in our loss term setting, we average the values over these common areas. These parts are usually synthesized many times due to our selection of bootstrapping scenes in a single time. This ensures elements across different bootstrapping scenarios are normalized.
@@
During the densification process, where points are either split or cloned, the surrounding points of these imperfect sections are initially flattened to compensate for the inconsistencies in the regenerated novel-view renderings. After several iterations’ accumulation, big points are split into small ones, and small points are cloned along the gradient direction. Our primary focus is on the cloning aspect since large points are unable to represent fine-grained details within a scene. In practical terms, these large points can be considered as coarse aggregations of smaller Gaussian distributions, where the renderings of these formations appear blurred. Over time, as the 3D-GS model progresses and refines, these big points gradually diminish and are refined into finer details. As such, they typically cease to exist in the final stages of the model’s training. Consequently, while they may be noticeable in earlier phases, their overall impact on the final output of the model is minimal.
The compensation process tends to be unstable, as the renderings vary even within the same scene section. However, as previously mentioned, we utilize multiple regenerated novel-view renderings to train a single view. For these renderings, if the details within them are not in conflict, then they represent unseen details or coarse modeling that we can modify freely. Otherwise, if there are conflicts among them, we can choose the most universally applicable ones to align. Given one point that suffers from different gradient directions to update, noted as $\sum_{i\in m}{\partial g_{m}}$, where m is the total number of different gradient directions and $\partial g_{m}$ the corresponding gradient. If we abstract these high-dimensional gradient directions into simple two-dimensional directions, we can further simplify this problem.
@@
Assuming each gradient in a 2D xy-plane comprises only two directions, with positive y pointing toward the ground truth and x representing deviations for simplicity, we separate each direction into an accumulation form. We hypothesize that the diffusion model generally produces images aligned with the context but varies in detail. Then by establishing an appropriate threshold for gradient accumulation magnitude and integrating it with our optimization strategy outlined in Eq. 8, we can achieve the desired optimization outcome, as shown in Fig 3. After several rounds of splitting, points will be small enough and only a small amount of gradients will occur at those points (even if not, the $\lambda_{\text{boot}}$ is small enough). For further cloning, only these gradients that toward exactly one direction with small deviations have the potential to exceed that threshold, and then we get a relatively faithful bootstrapping cloning point that constitutes the detail in that scene.
@@ 4.1 Experimental Setup
Our experiments are conducted on publicly available mainstream 3D reconstruction datasets, following the 3D-GS [1]. They include seven scenes from Mip-NeRF360 [83], two scenes from Tanks&Temples [84], two scenes from DeepBlending [85]. The evaluation metrics are PSNR, SSIM [52], and LPIPS [53]), together with training time. Frame rate (FPS) and memory usage are not primary concerns in our study, given that our technique centers on bootstrapping. The diffusion process and the bootstrapping loss described in Eq. 8, inherently require additional time to execute. In the main paper, we present averaged metrics for all scenes from each dataset, while detailed quantitative results for individual scenes are provided in the supplementary materials.
@@
In the initial stage of our experiment, we perform bootstrapping at specific intervals, dedicating 1000 iterations to each bootstrapping span. During each of these 1000-iteration intervals, we set the $\lambda_{\text{boot}}$ to 0.15 for the first 500 iterations and reduce it to 0.05 for the remaining 500 iterations. In configuring the diffusion model, we employ a progressively decreasing image breakage strength $s_{r}$, ranging from 0.05 to 0.01, across 100 time steps of DDIM [28] sampling. For various datasets, we implemented different strategies for creating novel-view cameras (Random or Consecutive). The bootstrapping regenerated images are executed using the original SD2.1v model. For the second stage experiment, the SD2.1v model is finetuned separately on each training dataset while leaving other configurations unchanged. The decision to finetune the SD2.1v model was based on empirical findings from our initial stage experiments, not a mere whim. After completing these experiments, we observed significant fluctuations in our results: while some scenes showed clear progress in metrics, others exhibited noticeable degradation, which is further discussed in Sec. 4.2. We also provide a detailed explanation of this configuration and the finetuning strategy of SD2.1v model in our supplementary material.

### v2

@@ 2.1 Diffusion Models
where $\epsilon_{\theta}$ is a learned noise prediction function, $\bar{\alpha}_{t}$ the schedule factor, and $\epsilon$ a random normal. For more details, please see in [27, 28, 29]
@@
Image-to-image generation represents a practical application of diffusion models. This paper specifically addresses a unique form of generation wherein an input image is regenerated by the diffusion model, utilizing a defined broken strength $s_{r}$ to facilitate controlled generation. In this approach, only a partial reverse process is engaged. The magnitude of $s_{r}$ directly influences the degree of deviation in the output image from the original, with larger $s_{r}$ values resulting in greater alterations. For example, if the $s_{r}=0.1$ and the total time step $T=1000$, then during the reverse diffusion process, diffusion is directly applied at the time step $T\times s_{r}=100$. In other words, only the last 100 time steps of the reverse diffusion process are executed to conduct moderate regeneration.
@@ 3 Methods
Figure 3: Training Pipeline. For novel-view rendering with artifacts (a), we initially prepare multiple adjacent novel-view renderings, ensuring the inclusion of the imperfect segment in each view (b). Subsequently, we regenerate all the views utilizing the diffusion model (c). Finally, through the application of our tailored training strategies, the final result demonstrates significant enhancement in the quality of the imperfect part (d).
3.1 Boostrapping by Diffusion Model
For a novel-view image synthesized by the trained 3D-GS model, its fidelity to the ground truth image may be compromised due to the presence of unseen parts not included in the training dataset. In instances where this novel-view image is considered a ”partially degenerated” version of the ground truth, a diffusion model can be utilized to perform image-to-image integration on the synthesized image to approximate the ground truth image, as described in Sec. 2.1. Given novel-view RBG image $\mathbf{I}\in\mathbb{R}^{w\times h\times 3}$, there exists a differential $\delta_{i}\in\mathbb{R}^{w\times h\times 3}$ generated by diffusion models that facilitates this image integration:
$\mathbf{I}^{\prime}=\delta_{i}+\mathbf{I},$ (7)
where $w$ and $h$ are the width and height of $\mathbf{I}$, and $\mathbf{I}^{\prime}$ is the regenerated image. Then using the diffusion process, we can reformulate the $\delta_{i}$ as a collection of diffusion noise generated by the noise predictor:
$\delta_{i}=\sum_{t\in T_{s_{r}}}{{\boldsymbol{\epsilon}}_{\theta}(\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{t}(\mathbf{I},{\boldsymbol{\epsilon}})+\sqrt{1-\bar{\alpha}_{t}}{\boldsymbol{\epsilon}},t)},$ (8)
Here, $\mathbf{x}_{t}$ is a function of $\mathbf{I}$ and ${\boldsymbol{\epsilon}}$ (we write in this form just for simplicity and better comprehension, for detail please see [27, 28]).
@@ 3.1 Boostrapping by Diffusion Model
Under this framework, the degree to which the generated image variance $\delta_{i}$ aligns with the ground truth image variance $\delta_{i}^{\prime}$ directly correlates to the improvement in our results. However, achieving the ideal situation in practice is very difficult. This is not only because (1) the currently available open-source diffusion models cannot perform well in generating partial images in all scenarios, but also because (2) many scenarios themselves are complex and unique. Moreover, (3) the very process of bootstrapping introduces a level of uncertainty. Ideally, the constructed 3D scene would appear consistent across all perspectives. However, the variability introduced by the diffusion model makes it challenging to guarantee that scenes regenerated by the diffusion model from one new viewpoint will seamlessly align with those from another. Additionally, (4) not all parts of the synthesized novel-view images should be regenerated, if we apply the diffusion process to the whole image, some already well-trained parts will be distorted to some extent. Therefore, based on these problems, we have explored a more universally usable and more stable method.
3.2 Bootstrapping Design
Bootstrapping Issues.
Regarding the issue of overall scene consistency, it requires the start of bootstrapping from a relatively complete stage on reconstructed 3D scenes with small diffusion broken strength. In the context of image regeneration using diffusion models, if the surrounding context is homogeneous and exhibits similarity, it can be inferred that the regenerated content within our setting demonstrates minimal variation. This observation arises because the diffusion process is applied only during several final time steps. By definition, the scale factor $\sqrt{1-\bar{\alpha}_{t}}$, which governs the magnitude of noise added, approaches zero as the time schedule $t$ approaches zero, as delineated in [6, 27]. We constrain the range of the time schedule to ensure that the value of $\sqrt{1-\bar{\alpha}_{t}}$ remains small. This constraint guarantees that the value of $\delta_{i}$ is also maintained at a minimal level.
@@ 3.2 Bootstrapping Design
We bootstrap every training image with its novel-view variants. To get a novel-view image $\mathbf{I}_{t}^{n}$ from a training image $\mathbf{I}_{t}$, we first construct a set of new camera parameters for rendering. In the COLMAP [19, 20] form, we solely slightly change a training camera’s $qvec$ and $tvec$ to obtain new view cameras, as detailed further in Sec. 4.1. Then we regenerate $\mathbf{I}_{t}^{n}$ by diffusion model and get $\mathbf{I}_{t}^{b}$
Bootstrapping Loss.
@@
The regenerated image $\mathbf{I}_{t}^{b}$ can be categorized into two parts: one that contains a faithful representation of the scene but changed with alterations, and another that exhibits an incomplete or inaccurate depiction of the scene, having been modified by the diffusion model. By managing the ratio of the loss associated with the original training component and that of the bootstrapping component, we can achieve a globally stable training process for the 3D-GS model and attain superior results.
@@
We set our bootstrapping loss term $\mathcal{L}_{b}$ as $\mathcal{L}_{1}$ loss of the novel-view rendered image $\mathbf{I}_{t}^{n}$ with the regenerated image $\mathbf{I}_{t}^{d}$, where $\mathcal{L}_{b}=\lVert\mathbf{I}_{t}^{n}-\mathbf{I}_{t}^{d}\lVert$. During training, if bootstrapping is used, then for every training camera, we use a hybrid loss with a small scaling factor $\lambda_{\text{boot}}$ that
@@
where $\mathcal{L}_{o}$ is the original 3D-GS training loss with ground-truth image [1], $\sum_{i\in N}{\mathcal{L}^{i}_{b}}$ is the cumulative bootstrapping loss across all bootstrapping views used in training this camera, and $N$ is the total variant number. The bootstrapping views employed are derived from all novel-view camera variants, which are based on the training camera and its surrounding cameras. For example, if the training camera has both sides of surrounding cameras and we bootstrap each camera with 2 of its variants, then the $N$ is 6. This design choice is motivated by the common characteristic of most 3D reconstruction datasets [83, 84, 85], where cameras positioned in close proximity typically capture adjacent scenes.
@@
When the $\lambda_{\text{boot}}$ is sufficiently small, such as 0.05, it becomes evident that the bootstrapping loss term will have minimal effects in the first scenario, in which well-trained parts are subject to alteration. For those parts or viewing angles that are underrepresented in the training dataset, where renderings appear visibly disordered or lack detail, the bootstrapping term can effectively facilitate modifications that remain consistent across multiple viewpoints. In our methodology, the bootstrapping scenarios consistently incorporate some of the same degenerated parts within our loss term configuration. Consequently, these parts are processed repeatedly from multiple angles, and the values are averaged over these areas to ensure uniformity and improve the overall quality of the synthesized views.
3.3 Methodological Analysis
In this section, our discussion will focus exclusively on the imperfect parts of the regenerated image (as described in the second scenario above), since the well-trained components of the model are relatively insensitive to our bootstrapping technique and can be easily rectified.
Emphasis on Cloning Direction.
@@ 3.3 Methodological Analysis
In the optimization process of 3D-GS, starting from a sparse set of COLMAP points, 3D-GS model will end up with millions of small and dense 3D Gaussians. Our bootstrapping technique particularly emphasizes the cloning function in training. During the densification process, where points are either split or cloned, the surrounding points of these imperfect sections are initially flattened to compensate for the inconsistencies in the regenerated novel-view renderings. After several iterations’ accumulation, big points are split into smaller ones, and small points are cloned along the gradient direction. In practical terms, these large points can be considered as coarse aggregations of smaller Gaussian points, where the renderings of these formations appear blurred. Over time, as the 3D-GS model progresses and refines, these big points gradually diminish and are refined into finer details. As such, they typically cease to exist in the final stages of the model’s training.
In conclusion, while large points may be noticeable in earlier phases, their overall impact on the final output of the model is minimal. Therefore, our primary focus should shift towards the cloning aspect. Specifically, we emphasize the gradient direction within the cloning procedure to ensure that finer details are accurately replicated and preserved, enhancing the overall fidelity and granularity of the scene.
Theoretical Analysis
The training process with bootstrapping tends to be unstable, as the regenerated renderings vary from each other even within the same scene section. We parameterize a Gaussian point with its position $\mu\in\mathbb{R}^{3}$, opacity $\alpha\in\mathbb{R}$, covariance-related quaternion $q\in\mathbb{R}^{4}$, scaling $s\in\mathbb{R}^{3}$, and color $c\in\mathbb{R}^{3}$. Each of these parameters is optimized simultaneously yet separately, allowing us to treat their gradients as independent. For the purpose of this analysis, we will consider a generalized situation.
The gradients generated by the bootstrapping term originate from the regeneration capabilities of the diffusion model, which prompts our initial focus on the diffusion process. Ideally, a well-trained diffusion model, as defined by nonequilibrium-thermodynamics diffusion models in [6], can reconstruct the original image from random noise through an infinite sequence of reverse diffusion sampling. Achieving this ideal scenario necessitates two conditions: (1) the model’s inherent capability to accurately simulate complex scenarios, and (2) an unlimited number of reverse time steps. In our experimental framework, we conduct only incomplete reverse diffusion processes with extremely limited time steps. Therefore, we can assume that our model adequately fulfills the capability requirements. Following condition (2), after infinite reverse sampling, all generated content should be identical, given the same initial states. If variations still occur after regeneration within the same defective segment, these can be attributed to the limited time step schedule. To address this issue, we can manually integrate reverse sampling into our process.
On a Gaussian point, for a gradient $\partial g_{b}\in\mathbb{R}^{d}$ generated through the bootstrapping image and another gradient $\partial g\in\mathbb{R}^{d}$ derived from the ground-truth image, where d is the dimension. it is reasonable to consider $\partial g_{b}$ as a degenerated version of $\partial g$ influenced by insufficient sampling:
$\partial g=\delta_{b}+\partial g_{b},$ (10)
where $\delta_{b}$ is the gradient difference. By the formulation specified in our loss term Eq.9, a Gaussian point typically encounters multiple bootstrapping gradients within a single training camera session. The bootstrapping differential $\delta_{i}$ in Eq 8, is conceptualized as a collection of noise, and our loss term is designed to facilitate the aggregation of these diverse noise collections. We can reformulate the bootstrapping loss $\sum_{i\in N}{\mathcal{L}^{i}_{b}}$ part as
$\sum_{i\in N}{\mathcal{L}^{i}_{b}}\propto\sum_{i\in N}{\sum_{t\in T_{s_{r}}}{{\boldsymbol{\epsilon}}_{\theta}(\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{t}(\mathbf{I}_{b}^{i},{\boldsymbol{\epsilon}})+\sqrt{1-\bar{\alpha}_{t}}{\boldsymbol{\epsilon}},t)}},$ (11)
where $\mathbf{I}_{b}^{i}$ is the $i$th bootstrapping image. Given our focus on the imperfect segments, and considering that each segment’s surrounding context is similar and homogeneous as described in Sec 3.2, we can approximate each $\mathbf{I}_{b}^{i}$ within that segment as identical. Then, for each defective segment, we can rewrite Eq.11:
$\sum_{i\in N}{\mathcal{L}^{i}_{b}}\propto N\sum_{t\in T_{s_{r}}}{{\boldsymbol{\epsilon}}_{\theta}(\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{t}(\mathbf{I}_{b},{\boldsymbol{\epsilon}})+\sqrt{1-\bar{\alpha}_{t}}{\boldsymbol{\epsilon}},t)},$ (12)
As such, we integrate the defective segments with N times repeated reverse sampling of bootstrapping. This process inherently compensates for the issue of insufficient sampling by aggregation, thereby enhancing content generation to be more consistent.
Practical Cloning Design
@@
Experimentally, we hypothesize that the diffusion model typically generates images that are contextually aligned but vary in detail. However, even with multiple times of reverse sampling, some normalized variances remain significantly large when compared to ground-truth images. For simplification, we project the $d$-dimensional gradient onto a 2D xy-plane, with the positive y-axis directed towards $\partial g_{b}$ and other directions indicating deviations. By setting an appropriate threshold for the gradient accumulation magnitude during cloning, we can achieve a more stable optimization outcome, as shown in Fig 4. After several rounds of splitting, points will be small enough and only a small amount of gradients will occur at those points (even if not, the $\lambda_{\text{boot}}$ is small enough). For further cloning, only these gradients oriented precisely in one direction with small deviations have the potential to exceed that threshold. Then we get a relatively faithful bootstrapping cloning point that constitutes the detail in that scene.
@@ 4.1 Experimental Setup
Our experiments are conducted on publicly available mainstream 3D reconstruction datasets, following the 3D-GS [1]. They include seven scenes from Mip-NeRF360 [83], two scenes from Tanks&Temples [84], and two scenes from DeepBlending [85]. The evaluation metrics are PSNR, SSIM [52], and LPIPS [53]), together with training time. The efficiency issues (e.g. rendering speed and storage size) are not primary concerns in our study, given that our technique centers on bootstrapping. The diffusion process and the bootstrapping loss described in Eq. 9, inherently require additional time to execute. In the main paper, we present averaged metrics for all scenes from each dataset, while detailed quantitative results for individual scenes are provided in the supplementary materials.
@@
In the initial stage of our experiment, we perform bootstrapping at specific intervals, dedicating 1000 iterations to each bootstrapping span. During each of these 1000-iteration intervals, we set the $\lambda_{\text{boot}}$ to 0.15 for the first 500 iterations and reduce it to 0.05 for the remaining 500 iterations. In configuring the diffusion model, we employ a progressively decreasing image broken strength $s_{r}$, ranging from 0.05 to 0.01, across 100 time steps of DDIM [28] sampling. For various datasets, we implemented different strategies for creating novel-view cameras (Random or Consecutive). The bootstrapping regenerated images are executed using the original SD2.1v model. For the second stage experiment, the SD2.1v model is finetuned separately on each training dataset while leaving other configurations unchanged. The decision to finetune the SD2.1v model was based on empirical findings from our initial stage experiments, not a mere whim. After completing these experiments, we observed significant fluctuations in our results: while some scenes showed clear progress in metrics, others exhibited noticeable degradation, which is further discussed in Sec. 4.2. We also provide a detailed explanation of this configuration and the finetuning strategy of SD2.1v model in our supplementary material.